# Trio-ViT: Post-Training Quantization and Acceleration for Softmax-Free Efficient Vision Transformer

### v1

@@ IV-A2 Observations on Quantization of Lightweight MSAs
When quantizing MatMuls in Eq. (4) within lightweight MSAs to 8-bit, we encounter notably worse results, which manifest as a “Not-a-Number” (NaN) issue. We find that this issue primarily arises from the quantization of divisors/denominators. As depicted in Figs. 4 (a) and (b), the wide range of values within divisors results in a reduced quantization resolution for smaller values when adopting the uniform quantization. Nevertheless, smaller values within divisors exhibit greater sensitivity compared to larger values. For instance, rounding a divisor of to during quantization results in an absolute difference of , yet it only doubles the final division results. In contrast, approximating the divisor of to incurs a mere absolute difference but increases the final results by a factor of . Particularly, changing a divisor of to yields a negligible absolute difference, but it causes the final results larger. These examples distinctly underscore the inherent incompatibility of uniform quantization for divisors, especially those exhibiting a wide range of values.
@@ IV-B Channel-Wise Migration for DW’s Inputs
As introduced in Challenge # 2 in Sec. IV-A1, there exist extreme inter-channel variations in DW’s inputs, making the vanilla layer-wise quantization unsuitable. Fortunately, owing to the distinctive algorithmic characteristic of DWConvs, which process each input channel independently and thus eliminates the summations along channels, we can directly employ channel-wise quantization to assign individual scaling factors for each input channel. This approach effectively addresses the above challenge without compromising hardware efficiency. Nonetheless, despite its potential advantages in enhancing quantization accuracy, it significantly increases the number of scaling factors, posing another challenge for their optimizations via LSQ [32], which is widely adopted to optimize scaling factors in layer-wise quantized activations.
To address this limitation, we propose to adopt channel-wise migration on top of the layer-wise quantization for DW’s inputs. Specifically, as shown in Fig. 5, weights in DWConvs feature the same channel number as inputs, and each weight channel functions as an independent filter dedicated to processing its corresponding input channel. This arrangement enables us to assign distinct scaling factors to each weight channel. Thus, filter-wise quantization is essentially the channel-wise quantization for DW’s weights. Consequently, inter-channel variations of activations can be seamlessly transferred to weights through a mathematically equivalent transformation employing channel-wise migration factors :
where , , and are the output, input, weight, and migration factor for channel, respectively. donate the quantization function, and are the layer-wise and channel-wise scaling factors for and , respectively. This approach greatly facilitates activation quantization without hindering weight quantization. Note that weights can be pre-transformed before deployment to eliminate the on-chip computation. As for activations that depend on the input images during inference and thus cannot be pre-processed, we can fuse with to obtain a fused scaling factor in advance, thus avoiding the on-the-fly transformation:
@@ IV-C Filter-Wise Shifting for PW2’s Inputs
To eliminate the inter-channel asymmetries in PW2’s inputs, inspired by [33], we propose filter-wise shifting to pre-process PW2’s inputs before quantization. As expressed in Eq. (9), it subtracts each input channel by its channel-wise mean , thus obtaining the calibrated input centered around zero (see Fig. 6 (c)) and significantly reducing value ranges (see the comparison between Fig. 3 (f) and fig. 6 (d)). To accommodate the above filter-wise shifting for activations and keep the same functionality as the original PW, the original bias of the output channel needs to be updated to following Eq. (10), where donate the input channel number. This can be pre-computed to eliminate the on-chip processing.
@@ IV-D Log2 Quantization for Divisors in MSAs
As illustrated in Figs. 4 (a) and (b), log2 quantization will allocate more bins to smaller values and vice versa. This inherent characteristic aligns with the algorithmic property of divisors in MSAs, where small values exhibit higher quantization sensitivity as explained in Sec. IV-A2. Thus, to improve the quantization resolution of small values in divisors of MSAs, we advocate adopting log2 quantization as follows:
@@ V-A1 Design Challenge # 1: Various Operation Types
Design Choice # 1: Multipliers-Adder-Tree Architecture. Considering that PWConvs are the dominant type of operations, one natural thought is to build our dedicated accelerator with the Multipliers-Adder-Tree (MAT) architecture, which is a typical design to efficiently support PWConvs with channel parallelism [34, 35]. Specifically, as illustrated in Fig. 7 (a), each processing element (PE) lane in the MAT engine is responsible for the multiplication (via multipliers), summation (via the adder tree), and accumulation (via the accumulator) along the input channel dimension to generate each output pixel (for PWConvs) or partial sum (dubbed psum, for generic convolutions). Thus, the parallelism within the PE lane of the MAT engine is along the input channel dimension to facilitate psum reuse. Besides, inputs are broadcast to different PE lanes and multiplied with different weight filters to produce output/psum pixels from different output channels, thus the parallelism among PE lanes is along the output channel dimension to enhance input reuse.
@@
Limitations. Although the MAT architecture can efficiently process PWConvs as well as easily support generic convolutions and MatMuls (which can be treated as PWConvs with large batch sizes), it has limited flexibility when handling DWConvs. Firstly, as for DWConvs, only psums generated from the same sliding window can be summed and accumulated, thus the achievable parallelism of each PE lane in MAT is limited by kernel sizes of DWConvs. Secondly, DWConvs in EfficientViT feature various kernel sizes ( and ) and strides ( and ), resulting in different sizes of sliding windows and distinct overlap patterns between adjacent sliding windows when conducting convolutions. This necessitates extra line buffers and substantial memory management overheads to support the multiplication and summation functionality within PE lanes for generating consecutive output pixels [34]. Thirdly, as there is no input reuse opportunity in DWConvs, this will lead to a high memory bandwidth requirement and otherwise low PE utilization when accommodating the output channel parallelism among PE lanes in the MAT architecture.
@@
Design Choice # 2: Reconfigurable Architecture. To solve the above limitations, we can consider the reconfigurable architecture depicted in Fig. 7 (b), which incorporates multiple Reconfigurable Multiplier-ACcumulation units (R-MACs). (i) As depicted in Fig. 7 (c), when executing generic convolutions, PWConvs, and MatMuls, this architecture can be configured to operate in the down-forward accumulation mode to achieve the same functionality as the MAT architecture. This means that each PE lane supports the input channel parallelism to achieve psum reuse, while different PE lanes favor output channel parallelism to exploit the input reuse opportunity. (ii) As depicted in Fig. 7 (d), when executing DWConvs, R-MAC can be configured to run in the self-accumulation mode, thus partial sums within each sliding window can be temporally accumulated within each R-MAC, making this kind of design inherently supports DWConvs with various kernel sizes. Specifically, R-MACs within each PE lane can spatially compute multiple output pixels from different output channels to boost throughput, thus the parallelism is along the output channel here (see Fig. 7 (d) right).
@@
Besides, weights can be broadcast to all PE lanes and multiplied with input pixels from different sliding windows to generate consecutive output pixels from the same output channel (see Fig. 7 (d) top left). By doing this, we can reuse overlaps among adjacent sliding windows with only several auxiliary registers and easily support DWConvs with different strides. For example, as shown in Fig. 7 (d) bottom left, where we take computations of the DWConv with a stride of executed on R-MACs arranged in the same row from PE lanes as the example. Initially, a sequence of input pixels is transmitted to input shift registers, which is then moved forward by cycles. During each cycle, the first pixels in shift registers are independently multiplied with the broadcast weight , generating the psums for consecutive output pixels. After cycles, the computation progresses to the next row of the input feature map and the corresponding filter, maintaining the same computation mode. This process is repeated until all rows are processed, resulting in output pixels. As for computations of DWConvs with a stride of , overlaps among adjacent sliding windows are spaced instead of successive. Thus, odd-column-indexed input pixels within each row are initially transmitted to shift registers for processing, followed by the even-column-indexed pixels. Weights also need to be broadcast following the same “first odd, then even” rule to accommodate this modified computation scheme. Thereby, the parallelism among PE lanes in this architecture is along the output feature map to enhance weight reuse.
@@
Limitations. Despite its flexibility in supporting all types of operations in EfficientViT, there exist reconfigurable overheads. Firstly, the overheads in computational resources and buffers: each R-MAC needs a high-bit adder and a psum register to support the self-accumulation. Secondly, the overheads in control logic: extra multiplexers are required to simultaneously support both two accumulation patterns, i.e., self-accumulation and down-forward accumulation.
@@
Our Proposed Design: Hybrid Architecture. Considering the fact that: (i) PWConvs are the dominant operations in EfficientViT [12] and the MAT architecture can efficiently support them, as well as (ii) EfficientViT incorporates various operation types, especially DWConvs with various kernel sizes and strides, and the R-MAC design can flexibly support them, we propose a hybrid architecture for our dedicated accelerator to marry the best of both worlds. Specifically, it consists of a MAT engine to efficiently process generic convolutions, PWConvs, and MatMuls, and a R-MAC engine to effectively support the above three operation types and DWConvs, thus enhancing flexibility while maintaining hardware efficiency.
@@
Offered Opportunity: Inter-Layer Pipeline. Besides the efficiency and flexibility of our proposed hybrid architecture, it also offers an opportunity for the inter-layer pipeline, saving data access costs. Specifically, DWConvs are exclusively executed on the R-MAC engine in our hybrid accelerator and are sandwiched by two PWConvs in MBConvs of EfficientViT. Thus, when the R-MAC engine handles DWConvs, the resulting outputs can be immediately transmitted to the idle MAT engine and serve as inputs to participate in the computation of the subsequent PWConvs. This integration enables the computation of DWConvs and their following PWConvs to be fused, leading to enhanced hardware utilization and a reduction in data access costs from off-chip.
@@ V-A2 Design Challenge # 2: Distinct Computational Pattern of Attention
As expressed in Eq. (4), after query and key undergo ReLU, there are five remaining steps for producing the final attention map : (i) MatMuls between and value ; (ii) toke-wise summation of ; then (iii) MatMuls of and outputs from step i; (iv) matrix-vector multiplications of and outputs from step ii; and finally (v) divisions between outputs from step iii (divisors) and step iv (dividends). Thanks to our log2 quantization for divisors as introduced in Sec. IV-D, the costly divisions can be substituted by hardware-efficient bit-wise shifts. Thus, it is evident that, besides multiplications, there are also element-wise summations and bit-wise shifts involved in computations of lightweight attention in EfficientViT. These multiplication-free element-wise operations are inherently incompatible with our multiplication-based PE arrays. Furthermore, they also exhibit low computational intensity, yielding increased bandwidth requirements or potential delays [36].
@@ V-B Micro Architecture
As shown in Fig. 8, our dedicated accelerator consists of computing cores and several global buffers (buffer B/C and output buffer). Each computing core includes several internal buffers (buffer A and auxiliary/divisor buffers) and multiple computing units (R-MAC engine, MAT engine, auxiliary processors, and re-/log2 quantization modules). Specifically, as for computing units in each computing core, the R-MAC engine comprises PE lanes, each containing R-MACs and can be reconfigured to operate in either self-accumulation or down-forward accumulation modes. This flexibility enables the effective processing of all multiplication-based operations in EfficientViT, including generic convolutions, DWConvs, PWConvs, and MatMuls, as introduced in Design Choice # 2 in Sec. V-A1. The MAT engine consists of PE lanes, each including multipliers, and is developed to efficiently handle multiple multiplication-based operations in EfficientViT, excluding DWConvs, as explained in Design Choice # 1 in Sec. V-A1. Besides, the log2 quantization module is developed to quantize divisors in Eq. (4) following steps outlined at the end of first paragraph in Sec. IV-D, thus boosting quantization accuracy as well as enabling the conversion of costly divisions into hardware-efficient bit-wise shifts, as discussed in Sec. IV-D. Our accelerator is also equipped with several low-cost auxiliary processors, such as the adder tree and shifter array, to accommodate the computation of multiplication-free operations (e.g., row-wise summation and bit-wise shift) in MSAs, as discussed in Our Proposed Solution in Sec. V-A2. Additionally, a re-quantization module is also needed to re-quantize outputs following Eq. (12), where and are quantized output , input, and weight, respectively, and are their corresponding scaling factors, and / are both positive integers. By doing this, the floating-point re-scaling factors are converted into dyadic numbers (DN), and the re-quantization process can be implemented by integer-only multiplications and bit-wise shifts [11, 8, 7], thus facilitating both intra- and inter-layer pipelines on-chip.
@@ V-C Inter- and Intra-Layer Pipelines
As introduced above, our dedicated accelerator incorporates both multiplication-based engines (R-MAC and MAT engines, with DWConvs limited to the former) and multiplication-free engines (auxiliary processors that are designed to expedite computations in MSAs). This architecture inherently offers opportunities for pipeline processing, where various operations can be simultaneously executed on distinct computing units, thereby enhancing hardware utilization and throughput. As for the inter-layer pipeline, as shown in Figs. 9 (a) and (b), when the R-MAC engine handles DWConv, the resulting outputs are first subtracted by the channel-wise mean obtained on the calibration data to implement the filter-wise shifting introduced in Sec. IV-C, aiming to facilitate activation quantization. After that, they undergo re-quantization through the re-quantization module before being stored in the auxiliary buffer. Then, they are promptly directed to the idle MAT engine to serve as inputs for subsequent PWConv computations. Given that DWConvs entail fewer computations compared to PWConvs, once the processing of the current DWConv is completed, the R-MAC engine can be reassigned to participate in the concurrent computation of PWConv, alongside the MAT engine.
Regarding intra-layer pipeline, (i) when the R-MAC engine processes for the head, are broadcast to the auxiliary adder tree to generate the vector via row-wise summations. This implies that steps i/ii in the Sec. V-A2 can be executed simultaneously on distinct computing units. Concurrently, (ii) the MAT engine sequentially preform multiplications between and 1) already obtained as well as 2) to generate divisors and dividends in Eq. (4) for the head, respectively. This means that steps iv/iii in the Sec. V-A2 are computed consecutively on the MAT engine. During this process, the firstly generated divisors are cached in the divisor buffer and then routed to the log2 quantization module for log2 quantization. (iii) Once the dividends are obtained, they can be re-quantized via the re-quantization module and then sent to the auxiliary shifter array along with the already log2-quantized divisors to conduct element-wise divisions via bit-wise shifts, thus obtaining the final outputs of MSA. Note that once the computation for of all heads is finished, the R-MAC engine can be reused to compute divisors and dividends, together with the MAT engine.

### v3

@@ IV-A2 Observations on Quantization of Lightweight MSAs
When quantizing MatMuls in Eq. (4) within lightweight MSAs to 8-bit, we encounter notably worse results, which manifest as a “Not-a-Number” (NaN) issue. We find that this issue primarily arises from the quantization of divisors/denominators in Eq. (4). As depicted in Figs. 4 (a) and (b), the wide range of values within divisors results in a reduced quantization resolution for smaller values when adopting the uniform quantization. Nevertheless, smaller values within divisors exhibit much greater sensitivity compared to larger values. For instance, rounding a divisor of to during quantization results in an absolute difference of , yet it only doubles the final division results. In contrast, approximating a divisor of to yields a negligible absolute difference, but it causes the final results larger. These examples distinctly underscore the inherent incompatibility of uniform quantization for divisors, especially those exhibiting a wide range of values.
@@ IV-B Channel-Wise Migration for DW’s Inputs
As introduced in Challenge # 1 in Sec. IV-A1, there exist extreme inter-channel variations in DW’s inputs, making the vanilla layer-wise quantization unsuitable. Fortunately, owing to the distinctive algorithmic characteristic of DW, which processes each input channel independently and thus eliminates the summations along different channels, we can directly employ channel-wise quantization to assign individual scaling factors for each input channel. This approach effectively addresses the above challenge without compromising hardware efficiency. Nonetheless, despite its potential advantages in enhancing quantization accuracy, it significantly increases the number of scaling factors for activations, posing another challenge for their optimizations via LSQ [33], which is widely adopted to optimize activation quantization.
To address this limitation, we propose to adopt channel-wise migration on top of the layer-wise quantization for DW’s inputs. Specifically, as shown in Fig. 5, weights in DW feature the same channel number as inputs, and each weight channel functions as an independent filter dedicated to processing its corresponding input channel. This arrangement enables us to assign distinct scaling factors to each weight channel. Thus, as depicted in Fig. 5 (b), filter-wise quantization is essentially the channel-wise quantization for DW’s weights. Consequently, inter-channel variations of activations can be seamlessly transferred to weights through a mathematically equivalent transformation employing channel-wise migration factors :
where , , and are the output, input, weight, and migration factor for channel, respectively. donate the quantization function, and are the layer-wise and channel-wise scaling factors for and , respectively. This approach greatly facilitates activation quantization without increasing learnable scaling factors of activations and impeding weight quantization. Note that weights can be pre-transformed before deployment to eliminate the on-chip computation. As for activations that depend on the input images during inference and thus cannot be pre-processed, we can fuse with to obtain a fused scaling factor in advance, thus avoiding the on-the-fly transformation:
@@ IV-C Filter-Wise Shifting for PW2’s Inputs
To eliminate the inter-channel asymmetries in PW2’s inputs (introduced in Challenge # 2 of Sec. IV-A1), inspired by [34], we propose filter-wise shifting to pre-process PW2’s inputs before quantization. As expressed in Eq. (9), we subtract each input channel by its channel-wise mean , thus obtaining the calibrated input centered around zero (see Fig. 6 (c)) and significantly reducing value ranges (see the comparison between Figs. 3 (f) and 6 (d)). To accommodate the above filter-wise shifting for activations and keep the same functionality as the original PW, the original bias of the output channel needs to be updated to following Eq. (10), where donate the input channel number. This bias update can be pre-computed to eliminate the on-chip processing.
@@ IV-D Log2 Quantization for Divisors in MSAs
As illustrated in Figs. 4 (a) and (b), log2 quantization will allocate more quantization bins to smaller values and vice versa. This inherent characteristic aligns with the algorithmic property of divisors in MSAs, where small values exhibit higher quantization sensitivity as explained in Sec. IV-A2. Thus, to improve the quantization resolution of small values in divisors of MSAs, we advocate adopting log2 quantization:
@@ V-A1 Design Challenge # 1: Various Operation Types
Design Choice # 1: Multipliers-Adder-Tree Architecture. Considering that PWConvs are the dominant type of operations, one natural thought is to adopt the Multipliers-Adder-Tree (MAT) architecture, which is a typical design to efficiently support PWConvs with channel parallelism [35, 36]. Specifically, as illustrated in Fig. 7 (a), each processing element (PE) lane in the MAT engine is responsible for the multiplication (via multipliers), summation (via the adder tree), and accumulation (via the accumulator) along the input channel dimension to generate each output pixel (for PWConvs) or partial sum (dubbed psum, for generic convolutions). Thus, the parallelism within the PE lane of the MAT engine is along the input channel dimension to facilitate psum reuse. Besides, inputs are broadcast to different PE lanes and multiplied with different weight filters to produce output/psum pixels from different output channels, thus the parallelism among PE lanes is along the output channel dimension to enhance input reuse.
@@
Limitations. Although the MAT architecture can efficiently process PWConvs as well as easily support generic convolutions and MatMuls (which can be treated as PWConvs with large batch sizes), it has limited flexibility when handling DWConvs. Firstly, as for DWConvs, only psums generated from the same sliding window can be summed and accumulated, thus the achievable parallelism of each PE lane in MAT is limited by kernel sizes of DWConvs. Secondly, DWConvs in EfficientViT feature various kernel sizes ( and ) and strides ( and ), resulting in different sizes of sliding windows and distinct overlap patterns between adjacent sliding windows when conducting convolutions. This necessitates extra line buffers and substantial memory management overheads to support the multiplication and summation functionality within PE lanes for generating consecutive output pixels [35]. Thirdly, the lack of input reuse opportunities within DWConvs will lead to either a high memory bandwidth requirement or low PE utilization when accommodating the output channel parallelism among PE lanes in the MAT architecture.
@@
Design Choice # 2: Reconfigurable Architecture. To solve the above limitations, the reconfigurable architecture depicted in Fig. 7 (b) can be considered, which incorporates multiple Reconfigurable Multiplier-ACcumulation units (R-MACs). (i) As depicted in Fig. 7 (c), when executing generic convolutions, PWConvs, and MatMuls, this architecture can be configured to operate in the down-forward accumulation mode to achieve the same functionality as the MAT architecture. This means that each PE lane supports the input channel parallelism to achieve psum reuse, while different PE lanes facilitate output channel parallelism to exploit input reuse. (ii) As depicted in Fig. 7 (d), when executing DWConvs, this architecture can be configured to run in the self-accumulation mode, thus partial sums within each sliding window can be temporally accumulated with each R-MAC, making it inherently supports DWConvs with various kernel sizes.
@@
Specifically, as illustrated Fig. 7 (d) right, different R-MACs within each PE lane can spatially compute multiple output pixels from different output channels, thus the parallelism is along the output channel here. Besides, as shown in Fig. 7 (d) top left, weights can be broadcast to all PE lanes and multiplied with input pixels from different sliding windows to generate consecutive output pixels from the same output channel. By doing this, we can reuse overlaps among adjacent sliding windows with only several auxiliary registers and easily support DWConvs with different strides. For example, as shown in Fig. 7 (d) bottom left, where we take computations of the DWConv with a stride of executed on R-MACs arranged in the same row from PE lanes as the example. Initially, a sequence of input pixels is transmitted to input shift registers, which is then moved forward by cycles. During each cycle, the first pixels in shift registers are independently multiplied with the broadcast weight , generating the psums for consecutive output pixels. After cycles, the computation moves to the next row of the input feature map and the corresponding filter, continuing in the same computation mode. This process is repeated until all rows are processed, resulting in output pixels. As for computations of DWConvs with a stride of , overlaps among adjacent sliding windows are spaced instead of successive. Thus, odd-column-indexed input pixels within each row are initially transmitted to shift registers for processing, followed by the even-column-indexed pixels. Weights also need to be broadcast following the same “first odd, then even” rule to accommodate this modified computation scheme. Thereby, the parallelism among PE lanes in this architecture is along the output feature map to enhance weight reuse.
@@
Limitations. Despite its flexibility in supporting all types of operations in EfficientViT, there exist reconfigurable overheads in two aspects. Firstly, the overheads in computational resources and buffers: each R-MAC needs a high-bit adder and a psum register to support the self-accumulation. Secondly, the overheads in control logic: extra multiplexers are required to simultaneously support both two accumulation patterns, i.e., self-accumulation and down-forward accumulation.
@@
Our Proposed Design: Hybrid Architecture. Considering the fact that: (i) PWConvs are the dominant operations in EfficientViT [13] and the MAT architecture can efficiently support them, as well as (ii) EfficientViT incorporates various operation types, especially DWConvs with various kernel sizes and strides, and the R-MAC design can flexibly support all of them, we propose a hybrid architecture for our dedicated accelerator to marry the best of both designs. Specifically, it consists of a MAT engine to efficiently process generic convolutions, PWConvs, and MatMuls, and a R-MAC engine to effectively support the above three operation types and DWConvs, thus enhancing flexibility while maintaining hardware efficiency.
@@
Offered Opportunity: Inter-Layer Pipeline. Besides the efficiency and flexibility of our proposed hybrid architecture, it offers an opportunity for inter-layer pipeline processing, thus saving data access costs. Specifically, DWConvs are exclusively executed on the R-MAC engine in our hybrid accelerator and are sandwiched by two PWConvs in MBConvs of EfficientViT. Thus, when the R-MAC engine handles DWConvs, the resulting outputs can be immediately transmitted to the idle MAT engine and serve as inputs to participate in the computation of the subsequent PWConvs. This integration enables the computation of DWConvs and their following PWConvs to be fused, leading to enhanced hardware utilization and reduced data access costs from off-chip.
@@ V-A2 Design Challenge # 2: Distinct Computational Pattern of Attention
As expressed in Eq. (4), after query and key undergo ReLU, there are five remaining steps for producing the final attention map : (i) MatMuls between and value ; (ii) toke-wise summation of ; then (iii) MatMuls of and outputs from step i; (iv) matrix-vector multiplications of and outputs from step ii; and finally (v) divisions between outputs from step iii (divisors) and step iv (dividends). Thanks to our log2 quantization for divisors as introduced in Sec. IV-D, the costly divisions can be substituted by hardware-efficient bit-wise shifts. Thus, it is evident that, besides multiplications, there are also element-wise summations and bit-wise shifts involved in computations of lightweight attention in EfficientViT. These multiplication-free element-wise operations are inherently incompatible with our multiplication-based PE arrays. Besides, they exhibit low computational intensity, yielding increased bandwidth requirements or potential delays [37].
@@ V-B Micro Architecture
As shown in Fig. 8, our dedicated accelerator consists of computing cores and several global buffers (buffer B/C and output buffer). Each computing core includes several internal buffers (buffer A and auxiliary/divisor buffers) and multiple computing units (R-MAC engine, MAT engine, auxiliary processors, and re-/log2 quantization modules). Specifically, as for computing units in each computing core, the R-MAC engine comprises PE lanes, each containing R-MACs and can be reconfigured to operate in either self-accumulation or down-forward accumulation modes. This flexibility enables the effective processing of all multiplication-based operations in EfficientViT, including generic convolutions, DWConvs, PWConvs, and MatMuls, as introduced in Design Choice # 2 in Sec. V-A1. The MAT engine consists of PE lanes, each including multipliers, and is developed to efficiently handle multiple multiplication-based operations in EfficientViT, excluding DWConvs, as explained in Design Choice # 1 in Sec. V-A1. Besides, the log2 quantization module is developed to quantize divisors in Eq. (4) following steps outlined at the end of the first paragraph in Sec. IV-D, thus boosting quantization accuracy as well as enabling the conversion of costly divisions into hardware-efficient bit-wise shifts. Our accelerator is also equipped with several low-cost auxiliary processors, such as the adder tree and shifter array, to accommodate the computation of multiplication-free operations (e.g., row-wise summation and bit-wise shift) in MSAs, as discussed in Our Proposed Solution in Sec. V-A2. Additionally, a re-quantization module is also incorporated to re-quantize outputs following Eq. (12), where and are quantized output , input, and weight, respectively, and are their corresponding scaling factors, and / are both positive integers. By doing this, the floating-point re-scaling factors are converted into dyadic numbers, allowing the re-quantization process to be implemented using integer-only multiplications and bit-wise shifts [12, 8, 7], thus facilitating both intra- and inter-layer pipelines on-chip.
@@ V-C Inter- and Intra-Layer Pipelines
As introduced above, our dedicated accelerator incorporates both multiplication-based engines (R-MAC and MAT engines, with DWConvs limited to the former) and multiplication-free engines (auxiliary processors that are designed to expedite computations in MSAs). This architecture inherently offers opportunities for pipeline processing, where various operations can be simultaneously executed on distinct computing units, thereby enhancing hardware utilization and throughput. As for the inter-layer pipeline, as shown in Figs. 9 (a) and (b), when the R-MAC engine handles DWConv, the resulting outputs are first subtracted by the channel-wise mean obtained on the calibration data to implement the filter-wise shifting introduced in Sec. IV-C, aiming to facilitate activation quantization. After that, they undergo re-quantization through the re-quantization module before being stored in the auxiliary buffer. Then, they are promptly directed to the idle MAT engine to serve as inputs for subsequent PWConv computations. Given that DWConvs entail much fewer computations compared to PWConvs, once the processing of the current DWConv is completed, the R-MAC engine can be reassigned to participate in the concurrent computation of PWConv, alongside the MAT engine.
Regarding the intra-layer pipeline, (i) when the R-MAC engine processes for the head, are broadcast to the auxiliary adder tree to generate the vector via row-wise summations. This implies that steps i/ii in the Sec. V-A2 can be executed simultaneously on distinct computing units. Concurrently, (ii) the MAT engine sequentially preform multiplications between and the already obtained as well as to generate divisors and dividends in Eq. (4) for the head, respectively. This means that steps iv/iii in the Sec. V-A2 are computed consecutively on the MAT engine. During this process, the firstly generated divisors are cached in the divisor buffer and then routed to the log2 quantization module for log2 quantization. (iii) Once the dividends are obtained, they can be re-quantized via the re-quantization module and then sent to the auxiliary shifter array along with the already log2-quantized divisors to conduct element-wise divisions via bit-wise shifts, thus obtaining the final outputs of MSA. Note that once the computation for of all heads is finished, the R-MAC engine can be reused to compute divisors and dividends, together with the MAT engine.